OpenStudy (anonymous):

y-2=3/4(x-0) This was pulled from the graph, and it is asking for the equation.

OpenStudy (rmrjr22):

y = mx+b

OpenStudy (anonymous):

Okay, what is the answer of y-2=3/4(x-0) if y=mx +b

OpenStudy (rmrjr22):

3/4x-0+2=y

OpenStudy (anonymous):

y-2=3/4(x-0) i.e. y-2=3/4(x-0) i.e. 4y-8=3x \[i.e. y=\frac{ 3 }{ 4 } x + 2\] Which is the required eq of line.

OpenStudy (rmrjr22):

Bingo!

OpenStudy (rmrjr22):

No best response?n :(